# Optimistic Safety for Linearly-Constrained Online Convex Optimization

### v1

@@ A.5 Completing the proof
We can apply Lemmas 4.2 and 4.3 and condition on $\mathcal{E}_{\mathrm{conf}}$ and $\mathcal{E}_{\mathrm{azuma}}$ to get that
@@
$=\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})+\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})$
@@
Also, conditioned on $\mathcal{E}_{\mathrm{conf}}$, it holds that $x_{t}\in\mathcal{Y}_{j_{t}}^{p}\subseteq\mathcal{Y}$ for all $t\in[T]$ and therefore $Ax_{t}\leq b$ for all $t\in[T]$. To get a bound of the probability of both events holding, we use Theorem 3.1 and Lemma A.1 with the union bound to get that
@@ A.6 Side lemmas
Lemma A.6 (Elliptic Potential, Lemma 11 in Abbasi-Yadkori et al. (2011)).
@@
Lemma A.7 (Number of Phases).
@@
Because $\det(\bar{V}_{j+1})\geq 2\det(\bar{V}_{j})$ for all $j\in[1,N-1]$ by definition, it holds that $\det(V_{T})\geq\det(\bar{V}_{N})\geq 2^{N-1}\det(\bar{V}_{1})$. Then, taking the log of both sides and using the fact that $N\leq 1+2\log(2^{N-1})$, we have
$N$ $\leq 1+2\log\left(\frac{\det(V_{T})}{\det(\bar{V}_{1})}\right)$
$=1+2\log\det(\bar{V}_{1}^{-1/2}V_{T}\bar{V}_{1}^{-1/2})$
$\leq 1+2d\log\|\bar{V}_{1}^{-1/2}V_{T}\bar{V}_{1}^{-1/2}\|$
$=1+2d\log\left\|\bar{V}_{1}^{-1/2}\left(\sum_{t=1}^{T}x_{t}x_{t}^{\top}+\bar{V}_{1}\right)\bar{V}_{1}^{-1/2}\right\|$
$=1+2d\log\left\|\frac{1}{\lambda}\sum_{t=1}^{T}x_{t}x_{t}^{\top}+I\right\|$
$\leq 1+2d\log\left(\frac{1}{\lambda}\sum_{t=1}^{T}\left\|x_{t}x_{t}^{\top}\right\|+1\right)$
$=1+2d\log\left(\frac{1}{\lambda}\sum_{t=1}^{T}\left\|x_{t}\right\|^{2}+1\right)$
$\leq 1+2d\log\left(\frac{TD^{2}}{\lambda}+1\right)$
$\leq 1+2d\log\left(T+1\right)$
$\leq 4d\log\left(T\right),$
where the second-to-last inequality uses $\lambda\geq D^{2}$ and the last inequality uses $T\geq 3$. ∎
Appendix B Proof of Proposition 5.1 (Zero Expected Violation)
In this section, we restate and prove Proposition 5.1.
Proposition B.1 (Restatement of Proposition 5.1).
Consider the same setup as Theorem 4.1 except with $b\leftarrow b-\mu\mathbf{1}$ for $\mu\in[0,b_{\mathrm{min}})$. Then, it holds that $\mathbb{E}[y_{t}]\leq b$ for all $t\in[T]$ and
$\mathbb{E}[R_{T}]\leq 8DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{2dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}+2DG\sqrt{2T\log(SDT/b_{\mathrm{min}})}+GD+\frac{b_{\mathrm{min}}G}{S}$
when $\mu=b_{\mathrm{min}}/T$ and $\delta=\min(1/2,b_{\mathrm{min}}/(2SDT))$.
@@
Let the event that Theorem 4.1 holds be denoted by $\mathcal{E}$ which occurs with probability at least $1-2\delta$. Then, running OSOCO with $b\leftrightarrow b-\mu\mathbf{1}$ ensures that $Ax_{t}\leq b-\mu\mathbf{1}$ for all $t\in[T]$ conditioned on $\mathcal{E}$. Therefore,
$\mathbb{E}[y_{t,i}]$ $=\mathbb{E}[a_{i}^{\top}x_{t}+\epsilon_{t,i}]$
$=\mathbb{E}[a_{i}^{\top}x_{t}]$
$=\mathbb{E}[a_{i}^{\top}x_{t}\mathbb{I}\{\mathcal{E}\}]+\mathbb{E}[a_{i}^{\top}x_{t}\mathbb{I}\{\mathcal{E}^{c}\}]$
$\leq b_{i}-\mu+2SD\delta$
@@
where we use the trivial constraint bound $a_{i}^{\top}x_{t}\leq SD$ in the first inequality and then the choice of $\mu$ and $\delta$ in the last step. This gives the safety guarantee.
Next, we study the impact on the regret. Let $\tilde{\mathcal{Y}}=\{x\in\mathcal{X}:Ax\leq b-\mathbf{1}\mu\}$. First, note that when $\mu\in[0,b_{\mathrm{min}}]$, it holds that $\alpha x^{\star}$ is in $\tilde{\mathcal{Y}}$ where $\alpha=\frac{b_{\mathrm{min}}-\mu}{b_{\mathrm{min}}}$ as
$a_{i}^{\top}(\alpha x^{\star})=\alpha a_{i}^{\top}x^{\star}\leq\alpha b_{i}=b_{i}\frac{b_{\mathrm{min}}-\mu}{b_{\mathrm{min}}}\leq b_{i}\frac{b_{i}-\mu}{b_{i}}=b_{i}-\mu$
for all $i\in[n]$, where we use the fact that the form $\frac{y-\mu}{y}$ is increasing in $y$ when $y>0$. Therefore, we know that
$\|\Pi_{\tilde{\mathcal{Y}}}(x^{\star})-x^{\star}\|\leq\|\alpha x^{\star}-x^{\star}\|=(1-\alpha)\|x^{\star}\|\leq\frac{\mu}{b_{\mathrm{min}}}D.$
Finally, with $\bar{x}^{\star}=\operatorname*{arg\,min}_{x\in\tilde{\mathcal{Y}}}\sum_{t=1}^{T}f_{t}(x)$ and $R_{T}^{\mathrm{osoco}}$ as the (high probability) regret bound of OSOCO, it holds conditioned on $\mathcal{E}$ that
@@
$\leq R_{T}^{\mathrm{osoco}}+\sum_{t=1}^{T}f_{t}(\Pi_{\tilde{\mathcal{Y}}}(x^{\star}))-\sum_{t=1}^{T}f_{t}(x^{\star})$
$\leq R_{T}^{\mathrm{osoco}}+G\sum_{t=1}^{T}\|\Pi_{\tilde{\mathcal{Y}}}(x^{\star})-x^{\star}\|$
$\leq R_{T}^{\mathrm{osoco}}+GDT\frac{\mu}{b_{\mathrm{min}}}$
$\leq R_{T}^{\mathrm{osoco}}+GD$
where we use the choice of $\mu$ in the last step. However, we care about the expected regret, so we get
@@
Appendix C Proof of Propositions 6.1 and 6.2 (Improved Efficiency)
We give proofs of Propositions 6.1 (Relaxed OSOCO) and 6.2 (HedgeDescent) in Sections C.1 and C.2, respectively.
C.1 Proof of Proposition 6.1 (Relaxed OSOCO)
We restate and prove Proposition 6.1 in the following.
Proposition C.1 (Duplicate of Proposition 6.1).
Consider the same setup as Theorem 4.1 except with OSOCO modified such that $\mathcal{Y}_{j}^{o}\leftrightarrow\{\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)\}_{k,\xi}$. Then, it holds with probability at least $1-2\delta$ that
$R_{T}\leq\mbox{ }$ $8DG\frac{1}{b_{\mathrm{min}}}\beta_{T}d\sqrt{2T\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}$
and $Ax_{t}\leq b$ for all $t\in[T]$.
Proof.
First note that, following the same process as Lemma A.2, it holds that $\gamma_{t}\geq 1-\frac{1}{b_{\mathrm{min}}}\sqrt{d}\bar{\beta}_{j_{t}}\|\bar{V}_{j_{t}}^{-1/2}x_{t}\|_{\infty}$. Then, using the equivalence of norms, it follows that
Following the rest of the proof of Theorem 4.1 with this bound gives the result. ∎
C.2 Proof of Proposition 6.2 (HedgeDescent)
We restate and prove Proposition 6.2 in the following.
Proposition C.2 (Duplicate of Proposition 6.2).
Consider an OCO setting with action set $\bar{\mathcal{X}}=\bigcup_{m\in[M]}\mathcal{X}_{m}$ where each $\mathcal{X}_{m}$ is convex and $\|\nabla f_{t}(x)\|\leq G$ and $\|x\|\leq D/2$ for all $x\in\bar{\mathcal{X}}$ and $t\in[T]$ (as in Definition 3.2). If HedgeDescent (Algorithm 2) is played in this setting with $\zeta_{t}=\frac{\sqrt{4\log(M)}}{GD\sqrt{t}}$ and $\eta_{t}=\frac{D}{G\sqrt{t}}$, then it holds that
@@
Note that this update is exactly the form of the classical Hedge algorithm with step size $\bar{\zeta}_{t}=\sqrt{4\log(M)}/\sqrt{t}$. As such, using the standard analysis of Hedge with time-dependent step size, e.g. Chernov & Zhdanov (2010), it holds that
@@
$\sum_{t=1}^{T}\mathbb{E}_{m_{t}\sim p_{t}}[f_{t}(x_{t}(m_{t}))]-\sum_{t=1}^{T}f_{t}(x^{\star})$ $=\sum_{t=1}^{T}\mathbb{E}_{m_{t}\sim p_{t}}[f_{t}(x_{t}(m_{t}))]-\sum_{t=1}^{T}f_{t}(x_{t}(m^{\star}))+\sum_{t=1}^{T}f_{t}(x_{t}(m^{\star}))-\sum_{t=1}^{T}f_{t}(x^{\star})$
@@
Appendix D Example OCO Algorithms

### v3

@@ B.5 Completing the proof
We can apply Lemmas 5 and 6 and condition on $\mathcal{E}_{\mathrm{conf}}$ and $\mathcal{E}_{\mathrm{azuma}}$ to get that
@@
$=\underbrace{\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})}_{\text{Term I}}+\underbrace{\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})}_{\text{Term II}}$
@@
Also, conditioned on $\mathcal{E}_{\mathrm{conf}}$, it holds that $x_{t}\in\mathcal{Y}_{j_{t}}^{p}\subseteq\mathcal{Y}:=\{x\in\mathcal{X}:Ax\leq b\}$ for all $t\in[T]$ and therefore $Ax_{t}\leq b$ for all $t\in[T]$. To get a bound of the probability of both events holding, we use Theorem 2 and Lemma 3 with the union bound to get that
@@ B.6 Side lemmas
Lemma 7 (Elliptic Potential, Lemma 11 in Abbasi-Yadkori et al. (2011)).
@@
Lemma 8 (Number of Phases).
@@
Because $\det(\bar{V}_{j+1})\geq 2\det(\bar{V}_{j})$ for all $j\in[1,N-1]$ by definition, it holds that
$\det(V_{T})\geq\det(\bar{V}_{N})\geq 2\det(\bar{V}_{N-1})\geq\cdots\geq 2^{N-1}\det(\bar{V}_{1})$
$\Longrightarrow\quad 2^{N-1}\leq\frac{\det(V_{T})}{\det(\bar{V}_{1})}$
$\Longleftrightarrow\quad(N-1)\log(2)\leq\log\left(\frac{\det(V_{T})}{\det(\bar{V}_{1})}\right)$
$\Longrightarrow\quad N\leq 1+2\log\left(\frac{\det(V_{T})}{\det(\bar{V}_{1})}\right)$
At the same time, it holds that,
$\frac{\det(V_{T})}{\det(\bar{V}_{1})}$ $=\det(\bar{V}_{1}^{-1/2}V_{T}\bar{V}_{1}^{-1/2})$
$\leq\|\bar{V}_{1}^{-1/2}V_{T}\bar{V}_{1}^{-1/2}\|^{d}$
$=\left\|\bar{V}_{1}^{-1/2}\left(\sum_{t=1}^{T}x_{t}x_{t}^{\top}+\bar{V}_{1}\right)\bar{V}_{1}^{-1/2}\right\|^{d}$
$=\left\|\frac{1}{\lambda}\sum_{t=1}^{T}x_{t}x_{t}^{\top}+I\right\|^{d}$
$\leq\left(\frac{1}{\lambda}\sum_{t=1}^{T}\left\|x_{t}x_{t}^{\top}\right\|+1\right)^{d}$
$=\left(\frac{1}{\lambda}\sum_{t=1}^{T}\left\|x_{t}\right\|^{2}+1\right)^{d}$
$\leq\left(\frac{TD^{2}}{\lambda}+1\right)^{d}.$
Putting everything together,
$N\leq 1+2\log\left(\frac{\det(V_{T})}{\det(\bar{V}_{1})}\right)\leq 1+2d\log\left(\frac{TD^{2}}{\lambda}+1\right)\leq 1+2d\log\left(T+1\right)\leq 4d\log\left(T\right)$
where we use $\lambda\geq D^{2}$ and $T\geq 3$. ∎
Appendix C Proof of Proposition 1 (Expectation Guarantees)
In this section, we restate and prove Proposition 1.
Proposition 3 (Restatement of Proposition 1).
Choose $\delta=\min(1/2,b_{\mathrm{min}}/(2SDT))$ and $\kappa=b_{\mathrm{min}}/T$, and set $\lambda$ and $\beta_{t}$ the same as in Thereom 1. Then, if $T\geq 3$, it holds that $\mathbb{E}[Ax_{t}]\leq b$ for all $t\in[T]$ and
$\mathbb{E}[R_{T}]\leq\frac{6DG}{b_{\mathrm{min}}}\beta_{T}\sqrt{2dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}+GD+\frac{b_{\mathrm{min}}G}{S}.$
@@ Appendix C Proof of Proposition 1 (Expectation Guarantees)
Let the event that Theorem 1 holds be denoted by $\mathcal{E}$ which occurs with probability at least $1-2\delta$. We individually show the safety guarantee and regret guarantee in the following sections.
Safety guarantee:
Running OSOCO (with $\kappa>0$) ensures that $Ax_{t}\leq b-\kappa\mathbf{1}$ for all $t\in[T]$ conditioned on $\mathcal{E}$. Therefore, it holds for all $i\in[n]$ that
$\mathbb{E}[a_{i}^{\top}x_{t}]$ $=\mathbb{E}[a_{i}^{\top}x_{t}\mathbb{I}\{\mathcal{E}\}]+\mathbb{E}[a_{i}^{\top}x_{t}\mathbb{I}\{\mathcal{E}^{c}\}]$
$\leq b_{i}-\kappa+2SD\delta$
@@
where we use the trivial constraint bound $a_{i}^{\top}x_{t}\leq SD$ in the first inequality and then the choice of $\kappa$ and $\delta$ in the last step. This gives the safety guarantee.
Regret guarantee:
Let $\tilde{\mathcal{Y}}=\{x\in\mathcal{X}:Ax\leq b-\mathbf{1}\kappa\}$. First, we show that when $\kappa\in[0,b_{\mathrm{min}}]$, it holds that $\alpha x^{\star}$ is in $\tilde{\mathcal{Y}}$ where $\alpha:=\frac{b_{\mathrm{min}}-\kappa}{b_{\mathrm{min}}}$. To do so, we need that both (a) $\alpha x^{\star}\in\mathcal{X}$ and (b) $a_{i}^{\top}(\alpha x^{\star})\leq b_{i}-\kappa$ for all $i\in[n]$. Point (a) holds because $\alpha\in[0,1]$, $\mathbf{0}\in\mathcal{X}$, $x^{\star}\in\mathcal{X}$ and $\mathcal{X}$ is convex, so
$\alpha x^{\star}=\alpha x^{\star}+(1-\alpha)\mathbf{0}\in\mathcal{X}.$
Point (b) holds because, for all $i\in[n]$,
$a_{i}^{\top}(\alpha x^{\star})=\alpha a_{i}^{\top}x^{\star}\leq\alpha b_{i}=\frac{b_{\mathrm{min}}-\kappa}{b_{\mathrm{min}}}b_{i}\leq\frac{b_{i}-\kappa}{b_{i}}b_{i}=b_{i}-\kappa,$
where the first inequality uses that $x^{\star}$ is feasible (i.e. $Ax^{\star}\leq b$) and the second inequality uses the fact that $\frac{b_{\mathrm{min}}-\kappa}{b_{\mathrm{min}}}$ is increasing in $b_{\mathrm{min}}$ and $b_{\mathrm{min}}\leq b_{i}$.
Then, with $\bar{x}^{\star}=\operatorname*{arg\,min}_{x\in\tilde{\mathcal{Y}}}\sum_{t=1}^{T}f_{t}(x)$ and $R_{T}^{\mathrm{osoco}}$ as the (high probability) regret bound of OSOCO when $\kappa>0$, it holds conditioned on $\mathcal{E}$ that
@@
$\leq R_{T}^{\mathrm{osoco}}+\sum_{t=1}^{T}f_{t}(\alpha x^{\star})-\sum_{t=1}^{T}f_{t}(x^{\star})$
$\leq R_{T}^{\mathrm{osoco}}+G\sum_{t=1}^{T}\|\alpha x^{\star}-x^{\star}\|$
$\leq R_{T}^{\mathrm{osoco}}+(1-\alpha)GDT$
$\leq R_{T}^{\mathrm{osoco}}+\frac{\kappa}{b_{\mathrm{min}}}GDT$
$\leq R_{T}^{\mathrm{osoco}}+GD,$
where the first inequality uses the regret bound on OSOCO given that it plays within $\tilde{\mathcal{Y}}$ and $\bar{x}^{\star}$ is the optimal action within $\tilde{\mathcal{Y}}$, the second inequality is due to the fact that $\alpha x^{\star}$ is in $\tilde{\mathcal{Y}}$ and therefore $\sum_{t=1}^{T}f_{t}(\bar{x}^{\star})=\min_{x\in\tilde{\mathcal{Y}}}\sum_{t=1}^{T}f_{t}(x)\leq\sum_{t=1}^{T}f_{t}(\alpha x^{\star})$, and the last inequality uses the choice of $\kappa$.
We can then apply this to the expected regret to get that
@@
Finally, we need to account for the choice of $\kappa$ in the regret bound $R_{T}^{\mathrm{osoco}}$. We do so by applying Theorem 1 with $b\leftarrow b-\kappa\mathbf{1}$ to get that
$R_{T}^{\mathrm{osoco}}$ $=4DG\frac{1}{b_{\mathrm{min}}-\kappa}\beta_{T}\sqrt{3dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{4dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}$
$=4DG\frac{1}{b_{\mathrm{min}}(1-1/T)}\beta_{T}\sqrt{3dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{4dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}$
$\leq 6DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{4dT\log(T)}+2DG\sqrt{2T\log(1/\delta)},$
where we use $T\geq 3$. ∎
Appendix D Proof of Propositions 2 (Improved Efficiency)
We first prove the regret guarantees of HedgeDescent in Section D.1 and then give the proof of Proposition 2 in Section D.2.
D.1 Regret of HedgeDescent
We give the regret guarantees of HedgeDescent (Algorithm 2) in the following proposition.
Proposition 4.
Consider the online optimization setting described in Definition 1 except with action set $\bar{\mathcal{X}}=\bigcup_{m\in[M]}\mathcal{X}_{m}$ where each $\mathcal{X}_{m}$ is convex. If HedgeDescent (Algorithm 2) is played in this setting with $\zeta_{t}=\sqrt{4\log(M)}/GD\sqrt{t}$ and $\eta_{t}=D/G\sqrt{t}$, then it holds that
@@ D.1 Regret of HedgeDescent
Note that this update is exactly the form of the classical Hedge algorithm with step size $\bar{\zeta}_{t}=\sqrt{4\log(M)}/\sqrt{t}$. As such, using the standard analysis of Hedge with time-dependent step size, e.g. Chernov and Zhdanov (2010), it holds that
@@
$\sum_{t=1}^{T}\mathbb{E}_{m_{t}\sim p_{t}}[f_{t}(x_{t}(m_{t}))]-\sum_{t=1}^{T}f_{t}(x^{\star})$
$=\sum_{t=1}^{T}\mathbb{E}_{m_{t}\sim p_{t}}[f_{t}(x_{t}(m_{t}))]-\sum_{t=1}^{T}f_{t}(x_{t}(m^{\star}))+\sum_{t=1}^{T}f_{t}(x_{t}(m^{\star}))-\sum_{t=1}^{T}f_{t}(x^{\star})$
@@
D.2 Completing the proof
In the following, we restate and prove Proposition 2.
Proposition 5 (Duplicate of Proposition 2).
If OSOCO is modified such that $\mathcal{Y}_{j}^{o}\leftarrow\{\tilde{\mathcal{Y}}_{j}^{o}(k,\xi)\}_{k,\xi}$ and HedgeDescent is used for $\mathcal{A}$, then the following hold:
  • (Static constraints) Choosing the algorithm parameters as in Theorem 1 ensures, with probability at least $1-2\delta$, that $Ax_{t}\leq b$ and $R_{T}\leq\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$.
  • (Time-varying constraints) Choosing the algorithm parameters as in Proposition 1 ensures that $\mathbb{E}[y_{t}]\leq b$ and $\mathbb{E}[R_{T}]\leq\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$.
Proof.
First note that, following the same process as Lemma 4, it holds that $\gamma_{t}\geq 1-\frac{1}{b_{\mathrm{min}}}\sqrt{d}\bar{\beta}_{j_{t}}\|\bar{V}_{j_{t}}^{-1/2}x_{t}\|_{\infty}$. Then, using the equivalence of norms, it follows that
Then, following the rest of the proof of Theorem 1 gives that, under the conditions of Theorem 1,
$R_{T}\leq\mbox{ }$ $8DG\frac{1}{b_{\mathrm{min}}}\beta_{T}d\sqrt{2T\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}+2DG\sqrt{2T\log(1/\delta)},$
and $Ax_{t}\leq b\ \forall t$ with probability at least $1-2\delta$. Then, following the proof of Proposition 1, it holds under the conditions of Proposition 1 that
$\mathbb{E}[R_{T}]\leq\mbox{ }$ $8DG\frac{1}{b_{\mathrm{min}}}\beta_{T}d\sqrt{2T\log\left(T\right)}+C_{\mathcal{A}}\sqrt{2dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}+GD+\frac{b_{\mathrm{min}}G}{S},$
and $\mathbb{E}[y_{t}]\leq b\ \forall t$. Finally, applying the regret guarantees of HedgeDescent by taking $C_{\mathcal{A}}=DG\sqrt{\log(2d)}+3DG$ gives the result. ∎
Appendix E Proof of Corollary 1 (Stochastic Constraints)
In this section, we restate and prove Corollary 1.
Corollary 2.
Suppose that the cost functions and action set satisfy Assumptions 1 and 2. Also, assume that the constraint function satisfies Assumption 6, and let $\rho^{2}=(G_{g}D+2F)^{2}$. Then playing OSOCO (Algorithm 1) with the algorithm parameters chosen as in Proposition 1 ensures that $\mathbb{E}[g_{t}(x_{t})]\leq 0$ for all $t\in[T]$ and $\mathbb{E}[R_{T}^{\mathrm{stoch}}]=\tilde{\mathcal{O}}((d+C_{\mathcal{A}}\sqrt{d})\sqrt{T})$. Furthermore, with modifications in Proposition 2, it holds that $\mathbb{E}[R_{T}^{\mathrm{stoch}}]=\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$.
Proof.
To show the claim, we need to show that the setting of OCO with stochastic constraints is subsumed by OCO with static constraints and no violation in expectation, and the guarantees will follow immediately from Proposition 1. We do so by taking $g_{t}(x_{t})=y_{t}$, $A:=\mathbb{E}[A_{t}]$, $b:=\mathbb{E}[b_{t}]$ and $\epsilon_{t}=(A_{t}-A)x_{t}-b_{t}+b$.
First, let $\mathcal{F}_{t}=\sigma(x_{1},\epsilon_{1},...,\epsilon_{t-1},x_{t})$. Also, we restate Assumption 6 in terms of $A_{t}$ and $b_{t}$:
  1. (i)
    $\|a_{t,i}\|\leq G_{g}$ for all $i\in[n]$,
  2. (ii)
    $\|b_{t}\|\leq F$,
  3. (iii)
    $b>0$,
  4. (iv)
    $b$ is known.
First, we show how the constraint $g_{t}(x_{t})\leq 0$ is subsumed by the model in Section 2. Indeed,
$y_{t}=g_{t}(x_{t})=A_{t}x_{t}-b_{t}=Ax_{t}-b+\underbrace{(A_{t}-A)x_{t}-b_{t}+b}_{\epsilon_{t}}.$
It remains to show that Assumption 4 holds, i.e. that $\mathbb{E}[\epsilon_{t,i}|\mathcal{F}_{t}]=0$ and $\mathbb{E}[\exp(\lambda\epsilon_{t,i})|\mathcal{F}_{t}]\leq\exp(\frac{\lambda^{2}\rho^{2}}{2})\ \forall\lambda\in\mathbb{R}$. Indeed,
$\mathbb{E}[\epsilon_{t,i}|\mathcal{F}_{t}]=\mathbb{E}[(a_{t,i}-a_{i})^{\top}x_{t}|\mathcal{F}_{t}]-\mathbb{E}[b_{t,i}|\mathcal{F}_{t}]+b_{i}=\mathbb{E}[(a_{t,i}-a_{i})^{\top}|\mathcal{F}_{t}]x_{t}-\mathbb{E}[b_{t,i}]+b_{i}=\mathbb{E}[(a_{t,i}-a_{i})^{\top}]x_{t}=0,$
where the second equality uses that $x_{t}$ is $\mathcal{F}_{t}$-measurable and that $b_{t}$ is independent of $\mathcal{F}_{t}$, and the third equality uses that $A_{t}$ is independent of $\mathcal{F}_{t}$. Also, it holds almost surely that,
$|\epsilon_{t,i}|$ $=|(a_{t,i}-a_{i})^{\top}x_{t}-b_{t,i}+b_{i}|$
$\leq(\|a_{t,i}\|+\|a_{i}\|)\|x_{t}\|+|b_{t,i}|+|b_{i}|$
$=(\|a_{t,i}\|+\|\mathbb{E}[a_{t,i}]\|)\|x_{t}\|+|b_{t,i}|+|\mathbb{E}[b_{t,i}]|$
$\leq(\|a_{t,i}\|+\mathbb{E}[\|a_{t,i}\|])\|x_{t}\|+|b_{t,i}|+\mathbb{E}[|b_{t,i}|]$
$\leq G_{g}D+2F.$
Therefore, it follows from the Hoeffding Lemma that,
$\mathbb{E}[\exp(\lambda\epsilon_{t,i})|\mathcal{F}_{t}]\leq\exp\left(\frac{\lambda^{2}(G_{g}D+2F)^{2}}{2}\right),$
and therefore we can take $\rho^{2}=(G_{g}D+2F)^{2}$. Therefore, the constraint $g_{t}(x_{t})\leq 0$ is subsumed by the model in Section 2.
Finally, note that,
$\mathbb{E}[g_{t}(x_{t})]=\mathbb{E}[Ax_{t}]-b+\mathbb{E}[\epsilon_{t}]=\mathbb{E}[Ax_{t}]-b+\mathbb{E}[\mathbb{E}[\epsilon_{t}|\mathcal{F}_{t}]]=\mathbb{E}[Ax_{t}]-b.$
Therefore, the in-expectation guarantees in Proposition 1, $\mathbb{E}[Ax_{t}]\leq b$ imply that $\mathbb{E}[g_{t}(x_{t})]\leq 0$. At the same time, it holds that $\bar{g}(x)=\mathbb{E}[A_{t}]x-\mathbb{E}[b_{t}]=Ax-b$ and therefore the regret notions are identical, i.e. $R_{T}=R_{T}^{\mathrm{stoch}}$. ∎
Appendix F Example OCO Algorithms